OpenStudy (anonymous):

n - mk = mt^2, for m

OpenStudy (gw2011):

n-mk=mt^2 (I assume t is squared) n=mt^2+mk (rearranging this you get): mt^2+mk=n m(t^2+k)=n m=n/(t^2+k)

OpenStudy (anonymous):

Thanks!

OpenStudy (gw2011):

You're welcome